# ejcolrgb1,0,0 The MandelZoom project II: the impact of stellar feedback on black hole accretion through an -disc in dwarf galaxies with a resolved interstellar medium

### v1

@@ Abstract
Unveiling the origin of supermassive black holes (SMBHs) and their coevolution with galaxies remains a fundamental question in astrophysics (see e.g., seminal papers by Rees, 1984; Efstathiou & Rees, 1988; Kormendy & Richstone, 1995; Silk & Rees, 1998; King, 2003). Intermediate-mass black holes (IMBHs; –) may hold a key to understanding this process, unravelling if the coevolution holds all the way down to the dwarf galaxy regime, as well as providing us a unique insight into how this coevolutionary picture may have emerged at early cosmic times. A central challenge here is to determine how IMBHs may form, how abundant they are, and under which physical conditions they can efficiently accrete gas. Furthermore, clarifying when and how such accretion episodes make IMBHs observable, for example, through active galactic nucleus (AGN) activity, mergers with other compact objects, tidal disruption events (TDEs), or feedback, is essential for constraining black hole (BH)–galaxy coevolution (e.g., Wang & Merritt, 2004; Mezcua, 2017; Greene et al., 2020; Koudmani et al., 2022; Pucha et al., 2025).
@@
Several studies suggest that the growth of light seeds, originating from Pop III remnants (see e.g., Hirano et al., 2014), is generally inefficient within the shallow potential wells of their host mini-haloes and may be strongly suppressed by stellar and BH feedback (Johnson & Bromm, 2007; Smith et al., 2018b; Sassano et al., 2023), while their small masses make them prone to wandering (Pfister et al., 2019), potentially stranding them in off-centre orbits that prevent efficient gas accretion (for a review see e.g., Inayoshi et al., 2020).
@@
By contrast, more massive stellar systems such as nuclear star clusters (NSCs) offer particularly favourable cradles for IMBH formation and growth. NSCs are ubiquitous even in low-mass galaxies (Sánchez-Janssen et al., 2019; Neumayer et al., 2020; Hoyer et al., 2021; Poulain et al., 2025), though their formation pathways remain under active investigation (e.g., Fahrion et al., 2022a; Fahrion et al., 2022b; van Donkelaar et al., 2024; Lahén et al., 2025; Gray et al., 2025). Recent -body simulations further demonstrate that IMBHs can form within massive clusters in only a few to several tens of Myr, particularly under low-metallicity conditions (e.g., Katz et al., 2015; Fujii et al., 2024; Rantala et al., 2024; Rantala et al., 2025). These results highlight several physical mechanisms that make NSCs uniquely conducive to BH growth: (i) mass segregation and dynamical friction drive hierarchical BH mergers; (ii) their deep potentials, in the case of massive clusters, help BH retention against dynamical ejections from two- and three-body interactions as well as gravitational-wave recoil kicks (e.g., Rantala et al., 2024; Rantala et al., 2025); (iii) their dense stellar backgrounds fuel BHs through TDEs (see e.g., Lee et al., 2023; Rantala et al., 2024; Rantala et al., 2025; Chang et al., 2025); (iv) they channel gas inflows (Partmann et al., 2025) and promote circumnuclear disc (CND) formation (Shin et al., 2025); and (v) the NSC, absorbing angular momentum exchange with the CND, facilitates the circularisation of inflowing gas into an accretion disc and thereby promotes the growth of BH mass and spin (Shin et al., 2025).
@@
One of the primary stages where such complex gas–star–BH interactions unfold is the CND. On scales of tens of parsecs, CNDs are found within NSCs and act as reservoirs of gas surrounding the central BH, having long been a major focus of both observational and theoretical studies. The CND around Sgr A* has been extensively investigated over the past decades (Becklin et al., 1982; Genzel et al., 1985; Guesten et al., 1987; Sutton et al., 1990, and for the review, Bryant & Krabbe (2021)), and intense nuclear starbursts (Davies et al., 2007; Hicks et al., 2013) and gravitational instability of CNDs (Izumi et al., 2023) have been observed in nearby Seyfert galaxies. Theoretical and numerical efforts have also sought to elucidate the role of CNDs in the life cycles of galactic nuclei. Hydrodynamic simulations show that CNDs are gravitationally unstable, leading to the formation of clumps and stars, followed by partial dispersal driven by stellar feedback (Roškar et al., 2015; Schartmann et al., 2018; Solanki et al., 2023; Barna et al., 2025; Shin et al., 2025). Under starburst conditions, SN feedback and turbulence can transform a thin, rotation-supported disc into a turbulent, clumpy, geometrically thick torus, which contributes significantly to nuclear obscuration and facilitates gas inflow toward the SMBH (Wada & Norman, 2002; Wada et al., 2009; Dinh et al., 2021; Shin et al., 2025).
Despite significant progress, the formation mechanism and long-term fate of the CND remain highly debated. Some studies argue that the CND is a transient structure, repeatedly disrupted by energetic processes such as SN explosions, powerful stellar winds from OB and WR stars, AGN jets, and turbulence driven by toroidal magnetic fields (Mezger et al., 1989; Morris & Serabyn, 1996; Genzel et al., 2010; Requena-Torres et al., 2012; Lau et al., 2013; Hsieh et al., 2018; Wada et al., 2023; Solanki et al., 2023). In contrast, other works suggest that the CND can persist as a quasi-stable structure for up to 10 Myr (Vollmer & Duschl, 2001; Christopher et al., 2005; Dinh et al., 2021). The persistence of the CND is closely tied to large-scale galactic inflows and the interaction between the ISM and the nuclear environment, with observations of inflowing gas streams indicating that external gas supply plays a crucial role in sustaining the disc (e.g., Vollmer & Duschl, 2001; Liu et al., 2012; Hsieh et al., 2017). Understanding this complex interplay between star formation, stellar feedback, and CND evolution is therefore essential for constraining the growth of central BHs, and it is equally important to place these processes within the broader galactic context rather than considering the CND in isolation. However, simulations that simultaneously capture a realistic multiphase ISM, trace the long-term CND cycle, and resolve the self-gravitating accretion disc feeding the SMBH remain exceedingly rare (Hopkins et al., 2024; Shin et al., 2025).
@@ 2.2 Gravity and hydrodynamics
Radiative heating and cooling processes are modelled using the Grackle chemistry and cooling library111https://grackle.readthedocs.io/ (Smith et al., 2017), which tracks the non-equilibrium abundances of six primordial species: H i, H ii, He i, He ii, He iii, and free electrons. Metal-line cooling is included through tabulated rates computed with the photoionisation code Cloudy (Ferland et al., 2013). We also include a metagalactic UV background based on the Haardt & Madau (2012) model, with self-shielding applied according to the prescription of Rahmati et al. (2013), as implemented in Grackle.
@@ 2.3.1 Star formation: explicit IMF sampling
where is the free-fall time, and is the star formation efficiency, consistent with observed values in dense molecular regions (e.g., Krumholz & Tan, 2007). To ensure sufficient temporal resolution of the gas depletion process, we restrict the timestep of star-forming cells to be no larger than , in addition to other integration constraints. Cells with non-zero SFRs are then stochastically converted into collisionless star particles in proportion to their computed SFRs. Then we assign individual stellar masses to each newly formed star particle by explicitly sampling from the Kroupa (2001) initial mass function (IMF) over the range of –. These stars collectively represent the internal stellar content of the particle, and all feedback processes are directly coupled to the evolution of these individual stellar components222This approach differs from the conventional SSP approximation, in which each star particle is assumed to fully sample the IMF. Such simplified treatments fail to capture the small-scale spatio-temporal clustering of ionizing sources from OB stars, which significantly affects the ability of photoionisation feedback to regulate star formation in the ISM (see Smith, 2021).. For details on how we preserve the IMF shape while addressing discrepancies with the dynamical masses of individual star particles, we refer the reader to Smith (2021).
@@ 2.3.2 Early stellar feedback: photoionisation (PI) & photoelectric (PE) heating
As part of the early stellar (or pre-SN) feedback, we implement both photoionisation and photoelectric heating, sourced by the interstellar radiation field (ISRF) from massive stars. Feedback quantities are derived by interpolating lookup tables as a function of the zero-age main sequence (ZAMS) mass, based on the Kroupa (2001) IMF. To reduce complexity, the stellar metallicity is fixed at rather than interpolating across metallicity. Stellar lifetimes are taken from the PARSEC evolutionary tracks (Bressan et al., 2012), while binary evolution is neglected. Emission rates of far-UV (FUV; – eV) and extreme-UV (EUV or ionizing; eV) photons are drawn from the OSTAR2002 stellar atmosphere models (Lanz & Hubeny, 2003; Emerick et al., 2019). For simplicity, these photon outputs are assumed to remain fixed at their ZAMS values throughout each star’s lifetime.
@@ 2.3.3 SN feedback
Stars with ZAMS masses between and end their lives as core-collapse SNe, injecting mass, metals, energy, and momentum into the surrounding gas. Injection is applied to the host gas cell and its immediate neighbours, defined as cells sharing a face with the host. Feedback quantities are distributed to preserve isotropy. To account for unresolved momentum in the Sedov–Taylor phase of the SN remnant, we implement a mechanical feedback scheme; however, at the resolution used in this study, most SNe are sufficiently well resolved (Smith et al., 2021). Type Ia SNe, stellar winds, and runaway OB stars are not included in our model. Each SN is assigned a fixed explosion energy of erg, with progenitor-mass-dependent ejecta masses and metallicities taken from Chieffi & Limongi (2004).
@@ 2.4 Super-Lagrangian refinement around BHs
As in Paper I, we adopt the super-Lagrangian refinement strategy introduced by Curtis & Sijacki (2015) to accurately resolve the transport of mass and angular momentum from the ISM down to the self-gravity radius of the accretion disc and the BH system. In this scheme, a refinement region of radius is defined around the BH, within which the maximum allowed cell radius increases linearly with distance from the BH. Specifically, grows linearly from pc at the BH location to pc at the outer boundary of the refinement region, pc. Within this refinement volume, cells are split or merged such that their radii remain within a controlled range: , where the refinement factor regulates the permitted span in cell sizes. The innermost resolution limit, given by pc, is sufficient to resolve the self-gravitating radius of the accretion disc, whose radius always exceeds pc in our BH system. To prevent the artificial formation of star particles with extremely low masses in this high-resolution region, star formation is explicitly suppressed in gas cells with masses below .
@@ 2.5 Black hole accretion model
We follow the method introduced in Fiacconi et al. (2018) and Paper I, adopting a steady-state -accretion disc model to describe BH accretion. Our model measures the mass and angular momentum flux from the surrounding ISM at the self-gravity radius of the accretion disc. Following Fiacconi et al. (2018), we analytically compute the evolution of the accretion disc and the BH according to the Shakura & Sunyaev (1973) solution, incorporating the effect of Lense-Thirring precession on a viscous disc via the Bardeen & Petterson (1975) effect. For further details of the implementation, we refer the reader to Fiacconi et al. (2018) and Paper I, and provide a brief summary here.
@@
The BH and accretion disc are initialised with masses of and , respectively, and with the Eddington fraction of . The initial spin parameter of the BH is set to . These parameters uniquely determine the initial angular momenta of both the BH and the accretion disc. We initialise the angular momentum vectors of the BH and the disc to be aligned with that of the galactic disc.
@@ 3.2 Global SFRs
Next, the right panel of Fig. 3 compares how the SFR varies as a function of time when adopting different choices of star formation prescriptions, gas softenings, and initial gas metallicity compared to the full run. With the exception of the lowZ simulation, reassuringly, all runs exhibit broadly similar SFRs to the full run. In the lowZ case, the lower metallicity reduces cooling efficiency, resulting in a newly-formed stellar mass that is about % of that in the full run (see Table 1). All simulations also show a significant decline in SFR during the first Myr, which can be attributed to the photoionisation feedback, as discussed earlier. Among them, the eff100 run exhibits the most intense star formation activity in the early phase, as expected, but the SFR subsequently drops to a similar level as in the other simulations, reaching . These results indicate that, at this resolution (), star formation becomes self-regulated by feedback, largely independent of the adopted star formation efficiency.
@@ 3.3 Time evolution of the multi-phase ISM
In the SN run, the overall gas density is the lowest among all simulations. In particular, after the initial starburst, the gas density outside the CND radius drops below , and the galactic disc becomes nearly quenched. As shown earlier in Fig. 2, when only SN feedback is included, clustered SNe heat the ISM very efficiently, suppressing both gas cooling and inflow into the CND, while gas draining onto the -disc continues. As a result, the CND mass gradually decreases, and its density remains the lowest among all simulations. By Myr, however, the heated ISM cools again, leading to renewed star formation and gas inflow onto the CND.
@@
Next, we consider the lower panels, which show simulations that adopt the same stellar feedback combination as the full run while varying star formation, gas softening and initial gas metallicity choices. As expected, they show qualitatively similar behaviour to the full run: (1) episodic star formation in the central – pc region, (2) – inflow episodes over a time period of 450 Myr, and (3) the CND existing only episodically, with lifetimes ranging from as short as Myr to more than Myr. In most runs, star formation in the inner pc region is suppressed because, due to the super-Lagrangian refinement, the cell mass decreases with radius and only a few cells satisfy the Jeans criterion (). In the fixed-SFthr run, however, star formation is allowed under conditions similar to those outside the refinement region, enabling stars to form even within the inner 0.1 pc of the CND. Interestingly, despite this more generous star formation criterion, the cumulative stellar mass in the CND is not significantly larger, because the feedback from in-situ stars self-regulates star formation within the CND itself (see Table 1).
@@ 3.4.1 In-situ star formation in the CND versus BH accretion rate
Examining the and BH accretion rate in more detail in the absence of feedback, the simulations show only small-amplitude, rapid oscillations in both and BH accretion rate555These rapid oscillations reflect the BH wobbling around the centre of the CND within a radius of pc., except at and Myr when clump-CND mergers (see Fig. 4) induce angular momentum transport and temporarily enhance gas inflow. Interestingly, the mass inflow rate onto the accretion disc in the PIPE run is slightly higher than in the noFB run. Following each SFR peak, the BH accretion rate shows a brief but distinct rise within Myr, caused by feedback-driven compression of gas that enhances the density inside 666More precisely, feedback briefly expels a significant amount of gas from the central region near the BH. This heated gas then falls back under the influence of the NSC potential on a free-fall timescale of Myr, leading to an enhanced density in the region.. In addition, because early star formation is inefficient in the PIPE run, the CND retains more mass than in noFB. This larger reservoir, combined with the effects of stellar feedback, leads to an overall higher BH accretion rate.

### v2

# ejcolrgb0,0,0 The MandelZoom project II: the impact of stellar feedback on black hole accretion through an -disc in dwarf galaxies with a resolved interstellar medium

@@ Abstract
Unveiling the origin of supermassive black holes (SMBHs) and their coevolution with galaxies remains a fundamental question in astrophysics (see e.g., seminal papers by Rees, 1984; Efstathiou and Rees, 1988; Kormendy and Richstone, 1995; Silk and Rees, 1998; King, 2003). Intermediate-mass black holes (IMBHs; –) may hold a key to understanding this process, unravelling if the coevolution holds all the way down to the dwarf galaxy regime, as well as providing us a unique insight into how this coevolutionary picture may have emerged at early cosmic times. A central challenge here is to determine how IMBHs may form, how abundant they are, and under which physical conditions they can efficiently accrete gas. Furthermore, clarifying when and how such accretion episodes make IMBHs observable, for example, through active galactic nucleus (AGN) activity, mergers with other compact objects, tidal disruption events (TDEs), or feedback, is essential for constraining black hole (BH)–galaxy coevolution (e.g., Wang and Merritt, 2004; Mezcua, 2017; Greene et al., 2020; Koudmani et al., 2022; Pucha et al., 2025).
@@
Several studies suggest that the growth of light seeds, originating from Pop III remnants (see e.g., Hirano et al., 2014), is generally inefficient within the shallow potential wells of their host mini-haloes and may be strongly suppressed by stellar and BH feedback (Johnson and Bromm, 2007; Smith et al., 2018a; Sassano et al., 2023), while their small masses make them prone to wandering (Pfister et al., 2019), potentially stranding them in off-centre orbits that prevent efficient gas accretion (for a review see e.g., Inayoshi et al., 2020).
@@
By contrast, more massive stellar systems such as nuclear star clusters (NSCs) offer particularly favourable cradles for IMBH formation and growth. NSCs are ubiquitous even in low-mass galaxies (Sánchez-Janssen et al., 2019; Neumayer et al., 2020; Hoyer et al., 2021; Poulain et al., 2025), though their formation pathways remain under active investigation (e.g., Fahrion et al., 2022b, a; van Donkelaar et al., 2024; Lahén et al., 2025; Gray et al., 2025). Recent -body simulations further demonstrate that IMBHs can form within massive clusters in only a few to several tens of Myr, particularly under low-metallicity conditions (e.g., Katz et al., 2015; Fujii et al., 2024; Rantala et al., 2024, 2025). These results highlight several physical mechanisms that make NSCs uniquely conducive to BH growth: (i) mass segregation and dynamical friction drive hierarchical BH mergers; (ii) their deep potentials, in the case of massive clusters, help BH retention against dynamical ejections from two- and three-body interactions as well as gravitational-wave recoil kicks (e.g., Rantala et al., 2024, 2025); (iii) their dense stellar backgrounds fuel BHs through TDEs (see e.g., Lee et al., 2023; Rantala et al., 2024, 2025; Chang et al., 2025); (iv) they channel gas inflows (Partmann et al., 2025) and promote circumnuclear disc (CND) formation (Shin et al., 2025); and (v) the NSC, absorbing angular momentum exchange with the CND, facilitates the circularisation of inflowing gas into an accretion disc and thereby promotes the growth of BH mass and spin (Shin et al., 2025).
@@
One of the primary stages where such complex gas–star–BH interactions unfold is the CND. On scales of tens of parsecs, CNDs are found within NSCs and act as reservoirs of gas surrounding the central BH, having long been a major focus of both observational and theoretical studies. The CND around Sgr A* has been extensively investigated over the past decades (Becklin et al., 1982; Genzel et al., 1985; Guesten et al., 1987; Sutton et al., 1990, and for the review, Bryant and Krabbe (2021)), and intense nuclear starbursts (Davies et al., 2007; Hicks et al., 2013) and gravitational instability of CNDs (Izumi et al., 2023) have been observed in nearby Seyfert galaxies. Theoretical and numerical efforts have also sought to elucidate the role of CNDs in the life cycles of galactic nuclei. Hydrodynamic simulations show that CNDs are gravitationally unstable, leading to the formation of clumps and stars, followed by partial dispersal driven by stellar feedback (Roškar et al., 2015; Schartmann et al., 2018; Solanki et al., 2023; Barna et al., 2025; Shin et al., 2025). Under starburst conditions, SN feedback and turbulence can transform a thin, rotation-supported disc into a turbulent, clumpy, geometrically thick torus, which contributes significantly to nuclear obscuration and facilitates gas inflow toward the SMBH (Wada and Norman, 2002; Wada et al., 2009; Dinh et al., 2021; Shin et al., 2025).
Despite significant progress, the formation mechanism and long-term fate of the CND remain highly debated. Some studies argue that the CND is a transient structure, repeatedly disrupted by energetic processes such as SN explosions, powerful stellar winds from OB and WR stars, AGN jets, and turbulence driven by toroidal magnetic fields (Mezger et al., 1989; Morris and Serabyn, 1996; Genzel et al., 2010; Requena-Torres et al., 2012; Lau et al., 2013; Hsieh et al., 2018; Wada et al., 2023; Solanki et al., 2023). In contrast, other works suggest that the CND can persist as a quasi-stable structure for up to 10 Myr (Vollmer and Duschl, 2001; Christopher et al., 2005; Dinh et al., 2021). The persistence of the CND is closely tied to large-scale galactic inflows and the interaction between the ISM and the nuclear environment, with observations of inflowing gas streams indicating that external gas supply plays a crucial role in sustaining the disc (e.g., Vollmer and Duschl, 2001; Liu et al., 2012; Hsieh et al., 2017). Understanding this complex interplay between star formation, stellar feedback, and CND evolution is therefore essential for constraining the growth of central BHs, and it is equally important to place these processes within the broader galactic context rather than considering the CND in isolation. However, simulations that simultaneously capture a realistic multiphase ISM, trace the long-term CND cycle, and resolve the self-gravitating accretion disc feeding the SMBH remain exceedingly rare (Hopkins et al., 2024; Shin et al., 2025).
@@ 2.2 Gravity and hydrodynamics
Radiative heating and cooling processes are modelled using the Grackle chemistry and cooling library111https://grackle.readthedocs.io/ (Smith et al., 2017), which tracks the non-equilibrium abundances of six primordial species: H i, H ii, He i, He ii, He iii, and free electrons. Metal-line cooling is included through tabulated rates computed with the photoionisation code Cloudy (Ferland et al., 2013). We also include a metagalactic UV background based on the Haardt and Madau (2012) model, with self-shielding applied according to the prescription of Rahmati et al. (2013), as implemented in Grackle.
@@ 2.3.1 Star formation: explicit IMF sampling
where is the free-fall time, and is the star formation efficiency, consistent with observed values in dense molecular regions (e.g., Krumholz and Tan, 2007). To ensure sufficient temporal resolution of the gas depletion process, we restrict the timestep of star-forming cells to be no larger than , in addition to other integration constraints. Cells with non-zero SFRs are then stochastically converted into collisionless star particles in proportion to their computed SFRs. Then we assign individual stellar masses to each newly formed star particle by explicitly sampling from the Kroupa (2001) initial mass function (IMF) over the range of –. These stars collectively represent the internal stellar content of the particle, and all feedback processes are directly coupled to the evolution of these individual stellar components222This approach differs from the conventional SSP approximation, in which each star particle is assumed to fully sample the IMF. Such simplified treatments fail to capture the small-scale spatio-temporal clustering of ionizing sources from OB stars, which significantly affects the ability of photoionisation feedback to regulate star formation in the ISM (see Smith, 2021).. For details on how we preserve the IMF shape while addressing discrepancies with the dynamical masses of individual star particles, we refer the reader to Smith (2021).
@@ 2.3.2 Early stellar feedback: photoionisation (PI) & photoelectric (PE) heating
As part of the early stellar (or pre-SN) feedback, we implement both photoionisation and photoelectric heating, sourced by the interstellar radiation field (ISRF) from massive stars. Feedback quantities are derived by interpolating lookup tables as a function of the zero-age main sequence (ZAMS) mass, based on the Kroupa (2001) IMF. To reduce complexity, the stellar metallicity is fixed at rather than interpolating across metallicity. Stellar lifetimes are taken from the PARSEC evolutionary tracks (Bressan et al., 2012), while binary evolution is neglected. Emission rates of far-UV (FUV; – eV) and extreme-UV (EUV or ionizing; eV) photons are drawn from the OSTAR2002 stellar atmosphere models (Lanz and Hubeny, 2003; Emerick et al., 2019). For simplicity, these photon outputs are assumed to remain fixed at their ZAMS values throughout each star’s lifetime.
@@ 2.3.3 SN feedback
Stars with ZAMS masses between and end their lives as core-collapse SNe, injecting mass, metals, energy, and momentum into the surrounding gas. Injection is applied to the host gas cell and its immediate neighbours, defined as cells sharing a face with the host. Feedback quantities are distributed to preserve isotropy. To account for unresolved momentum in the Sedov–Taylor phase of the SN remnant, we implement a mechanical feedback scheme; however, at the resolution used in this study, most SNe are sufficiently well resolved (Smith et al., 2021). Type Ia SNe, stellar winds, and runaway OB stars are not included in our model. Each SN is assigned a fixed explosion energy of erg, with progenitor-mass-dependent ejecta masses and metallicities taken from Chieffi and Limongi (2004).
@@ 2.4 Super-Lagrangian refinement around BHs
[…]
@@ 2.5 Black hole accretion model
We follow the method introduced in Fiacconi et al. (2018) and Paper I, adopting a steady-state -accretion disc model to describe BH accretion. Our model measures the mass and angular momentum flux from the surrounding ISM at the self-gravity radius of the accretion disc. Following Fiacconi et al. (2018), we analytically compute the evolution of the accretion disc and the BH according to the Shakura and Sunyaev (1973) solution, incorporating the effect of Lense-Thirring precession on a viscous disc via the Bardeen and Petterson (1975) effect. For further details of the implementation, we refer the reader to Fiacconi et al. (2018) and Paper I, and provide a brief summary here.
@@
[…]
@@ 3.2 Global SFRs
[…]
@@ 3.3 Time evolution of the multi-phase ISM
[…]
@@
[…]
@@ 3.4.1 In-situ star formation in the CND versus BH accretion rate
Examining the and BH accretion rate in more detail in the absence of feedback, the simulations show only small-amplitude, rapid oscillations in both and BH accretion rate666These rapid oscillations reflect the BH wobbling around the centre of the CND within a radius of pc., except at and Myr when clump-CND mergers (see Fig. 4) induce angular momentum transport and temporarily enhance gas inflow. Interestingly, the mass inflow rate onto the accretion disc in the PIPE run is slightly higher than in the noFB run. Following each SFR peak, the BH accretion rate shows a brief but distinct rise within Myr, caused by feedback-driven compression of gas that enhances the density inside 777More precisely, feedback briefly expels a significant amount of gas from the central region near the BH. This heated gas then falls back under the influence of the NSC potential on a free-fall timescale of Myr, leading to an enhanced density in the region.. In addition, because early star formation is inefficient in the PIPE run, the CND retains more mass than in noFB. This larger reservoir, combined with the effects of stellar feedback, leads to an overall higher BH accretion rate.